OpenStudy (anonymous):

What is an equation for the line with slope 2/3 and y-intercept 9?

OpenStudy (anonymous):

y = (2/3)x + 9

OpenStudy (anonymous):

Well if you want it in slope intercept form: Starts with y= and then you put the slope with x, and then the y intercept. y=2/3 x +9

OpenStudy (mathstudent55):

The slope-intercept form of a line is y = mx + b, where m = slope, and b = y-intercept. Since you are given the slope and the y-intercept, all you need to do is substitute them into the slope-intercept form.